# Testing Hypotheses regarding Covariance and Correlation matrices with the R package CovCorTest

### v1

@@ Abstract
In addition to the commonly analyzed measures of location, dispersion measurements such as variance and correlation provide many valuable information. Consequently, they play a crucial role in multivariate statistics, which leads to tests regarding covariance and correlation matrices. Furthermore, also the structure of these matrices leads to important hypotheses of interest, since it contains substantial information about the underlying model. In fact, assumptions regarding the structures of covariance and correlation matrices are often fundamental in statistical modelling and testing. In this context, semi-parametric settings with minimal distributional assumptions and very general hypotheses are essential for enabling manifold usage. The free available package CovCorTest provides suitable tests addressing all aforementioned issues, using bootstrap and similar techniques to achieve good performance, particularly in small samples. Additionally, the package offers flexible specification options for the hypotheses under investigation in two central tests, accommodating users with varying levels of expertise, which results in high flexibility and user-friendliness at the same time. This paper also presents the application of CovCorTest for various issues, illustrated by multiple examples, where the tests are applied to a real-world dataset.
@@ Introduction
Tests for measure of location are widely used in the analysis of multivariate data, such as those arising from repeated measures designs or multiple measurements on the same subject. While expectation vectors are a classical choice for such analyses (Konietschke et al. (2020)), alternative parameters such as quantiles (Baumeister et al. (2024)), relative effects (Rubarth et al. (2022) or Dobler et al. (2020)) and others have been proposed. Withal, measures of dispersion contain substantial information and are, therefore, of central importance in multivariate settings. Especially covariance and correlation matrices are excellent parameters for analyses, which is, among other reasons, because:
@@
  • In statistical modeling, covariance and correlation matrices and their structures (e.g. autoregressive or compound symmetry) are essential for understanding and describing the underlying processes.
@@
Due to their importance, a variety of statistical procedures have been developed to test hypotheses concerning covariance matrices. Unfortunately, most existing methods rely on restrictive distributional assumptions, such as normality in Box (1953) and Zhong et al. (2017), or specific conditions on the characteristic function in Gupta and Xu (2006), which limits their applicability. Furthermore, several approaches are limited to specific settings, for example, two group settings as in Sakaori (2002) or Omelka and Pauly (2012). Moreover, they typically allow the investigation of only one or two concrete hypotheses, such as equality of covariance matrices (Yang and DeGruttola (2012) or Zhu et al. (2002)) or equality of two correlation matrices (Jennrich (1970)). This also holds for the software packages implementing these procedures, such as the R-packages psych(Revelle (2019)) or CovTestR(Barnard and Young (2018)).
@@
In contrast, the framework introduced in Sattler et al. (2022) allows a broad class of hypotheses, building on methodologies familiar from the analysis of location parameters. At the same time, the less restrictive semi-parametric setting used therein allows broad applicability. Thereto, test statistics are constructed as quadratic forms and Monte-Carlo techniques and resampling approaches are used to obtain the required quantiles and achieve good small sample performance. This approach was further expanded in Sattler and Pauly (2023) to quite general hypotheses regarding correlation matrices, where the focus was the Anova-type-statistic, and additionally, a Taylor-based Monte-Carlo approach was incorporated. The corresponding R package CovCorTest (Sattler et al. (2025)) includes the adequate tests from both papers to provide consistent procedures for investigating covariance and correlation matrices, and moreover also combined tests for equality of covariance and correlation matrices at the same time. Additionally, specific structures of covariance and correlation matrices - such as the Toeplitz form - are of great interest, as they allow for meaningful insights into the underlying model. Besides, specific structures are usual model assumptions, but tests regarding these are often insufficient. Most existing approaches are limited to specific structures such as compound symmetry (Winer et al. (1991) or Votaw (1948)) or first-order autoregressive (McKeown and Johnson (1996)) or focus exclusively on either covariance or correlation matrices (Steiger (1980)). Conversely, the recent work by Sattler and Dobler (2024) introduces a unified framework for testing a large class of structural hypotheses, known as the linear covariance model respectively the linear correlation model, and even further structures. These tests, which are related to the approaches of Sattler et al. (2022) and Sattler and Pauly (2023), were also integrated into the package to enable comprehensive and coherent testing of various aspects of covariance and correlation matrices in one place.
@@
test_correlation_structure and test_combined), which are as analogous as possible to make it more user-friendly. This uniform design allows researchers to apply the functions in analogous situations targeting different aspects of the data. Moreover, this approach allows for the investigation of all related hypotheses within a single package, while the coherent construction and methodology enhance the comparability of the results. Furthermore, to support users with varying levels of experience, the covariance and correlation tests allow not only the selection of predefined hypotheses, but also the specification of custom hypotheses by providing a hypothesis matrix and the corresponding vector. Since for the remaining tests, either only one hypothesis is possible (test_combined) or only predefined structures can be tested, no such differentiation is necessary there.
Because all underlying approaches and tests are based on empirical covariance and correlation matrices, our functions should only be used for metrical data, although further assumptions, like being a repeated measure design, are not required.
In the subsequent section, we will introduce the statistical multivariate model, the investigated null hypotheses, together with the underlying assumptions. Afterwards, we present the implemented inference procedures, followed by the applications from the R package CovCorTest illustrated by some examples using a real-world dataset. After all, a conclusion and outlook complete the paper.
@@
We consider a general multigroup setting with $a\in{\mathbb{N}}$ groups, where each group $i=1,...,a$ contains $n_{i}\in{\mathbb{N}}$ observations. The total number of observations is denoted by $N=\sum_{i=1}^{a}n_{i}$, allowing for potentially unbalanced designs in which group sizes may differ.
@@
[…]
@@
    $\operatorname{Var}(\boldsymbol{\epsilon}_{i1\ell})=\sigma_{i\ell}>0$,  $\ell=1,...,d$
@@
Since both covariance and correlation matrices are symmetric, it is natural to consider an appropriate vectorization to transform these matrices into vectors containing the same information as the original matrix. For the covariance matrices, the so-called half-vectorization operation $\operatorname{vech}$ is used, which extracts the upper-triangular (including diagonal) elements of a $d\times d$ matrix, yielding a $p=d(d+1)/2$ dimensional vector. The upper-half-vectorization $\operatorname{vech}^{-}$, on the contrary, only takes the values above the diagonal into a $p_{-}=d(d-1)/2$ dimensional vector, which is useful for correlation matrices whose diagonal entries are always one and hence contain no information. Now, based on hypothesis matrices $\boldsymbol{C}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}\times p}$ respectively $\boldsymbol{C}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}\times p_{-}}$ and vectors $\boldsymbol{\zeta}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}}$, $\boldsymbol{\zeta}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}}$ general linear hypotheses can be formulated. These hypotheses are defined based on the pooled covariance vector $\boldsymbol{v}=(\operatorname{vech}(\boldsymbol{V}_{1})^{\top},\dots,\operatorname{vech}(\boldsymbol{V}_{a})^{\top})^{\top}$ respectively the pooled correlation vector $\boldsymbol{r}=(\operatorname{vech}^{-}(\boldsymbol{R}_{1})^{\top},\dots,\operatorname{vech}^{-}(\boldsymbol{R}_{a})^{\top})^{\top}$ as
@@
For correlation matrices, analogous—but typically fewer—structures exist, which often arise as heterogeneous versions of other structures such as heterogenous autoregressive, heterogeneous compound symmetry or heterogenous Toeplitz.
Except for the autoregressive structure and all its versions, all of these covariance and correlation structures can be expressed within the general hypothesis framework using appropriate choices of $\boldsymbol{C}$ and $\boldsymbol{\zeta}$, as shown in Sattler and Dobler (2024). This highlights the versatility and generality of the presented model in capturing also a wide range of relevant covariance and correlation structures.
@@
To evaluate the hypotheses formulated in the previous section, we develop inference procedures based on the asymptotic distribution of vectorized covariance and correlation estimates.
For our hypotheses regarding covariance matrices we consider the vectorized empirical group covariances $\widehat{\boldsymbol{v}}_{i}:=\operatorname{vech}\left((n_{i}-1)^{-1}\sum_{i=1}^{n_{i}}(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})^{\top})\right)$ where $\overline{\boldsymbol{X}}_{i}$ is the sample mean of group $i$. Stacking these vectors across all $a$ groups yields the pooled vector $\widehat{\boldsymbol{v}}=(\widehat{\boldsymbol{v}}_{1}^{\top},...,\widehat{\boldsymbol{v}}_{a}^{\top})^{\top}$. Then, for developing a statistical test, we use that under assumption (A2) and (A3) for $i=1,\dots,a$, the group-wise covariance vectors satisfy
@@
In the same way, $\widehat{\boldsymbol{r}}_{i}$ denotes the empirical counterpart of the upper-half-vectorized correlation matrix of group $i$ and $\widehat{\boldsymbol{r}}$ is the corresponding pooled vector. Following the results in Sattler and Pauly (2023) (where also the exact form can be found), there exists a continuous matrix-valued function $\boldsymbol{M}\in{\mathbb{R}}^{p_{-}\times p}$ such that
$\sqrt{n_{i}}(\widehat{\boldsymbol{r}}_{i}-\boldsymbol{r}_{i})=\boldsymbol{M}(\boldsymbol{v}_{i},\boldsymbol{r}_{i})\cdot\sqrt{n_{i}}(\widehat{\boldsymbol{v}}_{i}-\boldsymbol{v}_{i})+\scriptstyle\mathcal{O}{}_{P}(1).$ (2)
@@
There exists also alternative versions of the ATS, which, for example, allow the approximation through an F-distribution (see e.g. Brunner et al. (2019) or Thiel et al. (2024)). However, for the statistics defined in (3) and (4), the asymptotic null distribution is known, enabling valid inference without stringent distributional assumptions.
@@
A Monte-Carlo technique is one useful approach to obtain the required quantiles of these weighted sums for this statistic. For the covariance matrix, with our consistent covariance estimator $\widehat{\boldsymbol{\Sigma}}$, we get estimators $\hat{\lambda}_{1},...,\hat{\lambda}_{m_{1}}$ for the unknown eigenvalues. With these weights we generate appropriate realisations $B_{1},...,B_{m_{2}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\chi_{1}^{2}$. These realizations and our estimators allow calculating a weighted sum $\sum\limits_{\ell=1}^{m_{2}}\hat{\lambda}_{\ell}B_{\ell}$, which converges in distribution against the above limit distribution. Repeating this procedure many times allows to calculate the required Monte-Carlo quantiles, and with an analogous procedure for the correlation for $\alpha\in(0,1)$, we get asymptotic level $\alpha$ tests through
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{MC}&=11\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{MC}\right),\\ &\varphi_{\boldsymbol{r}}^{MC}&=11\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{MC}\right).\end{array}$
Another technique for obtaining quantiles that requires more computational effort but usually yields better small-sample performance is a bootstrap approach. For hypotheses regarding covariance matrices, the parametric bootstrap mimics the limit distribution and afterwards the structure of the test statistic by generating bootstrap vectors $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}_{p}\left({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i}\right),$ based on given realizations $\boldsymbol{X}_{i1},...,\boldsymbol{X}_{in_{i}}$ with estimators $\widehat{\boldsymbol{\Sigma}}_{i}$. Then $\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$, the empirical covariance matrix of the bootstrap sample $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}$ and $\widehat{\boldsymbol{\Sigma}}^{\star}:=\bigoplus_{i=1}^{a}{N}/{n_{i}}\cdot\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$ are calculated. Following the bootstrap realization $ATS_{\boldsymbol{v}}^{\star}=N[\boldsymbol{C}\overline{\boldsymbol{Z}}^{\star}]^{\top}[\boldsymbol{C}\ \overline{\boldsymbol{Z}}^{\star}]\big{/}\operatorname{tr}(\boldsymbol{C}\widehat{\boldsymbol{\Sigma}}^{\star}\boldsymbol{C}^{\top})$ is calculated, using the bootstrap means $\overline{\boldsymbol{Z}}_{i}^{\star}=n_{i}^{-1}\sum_{k=1}^{n_{i}}\boldsymbol{Z}_{i}k^{\star}$ and $\overline{\boldsymbol{Z}}^{\star}=({\overline{\boldsymbol{Z}}_{1}^{\star}}^{\top},...,{\overline{\boldsymbol{Z}}_{a}^{\star}}^{\top})^{\top}$. Then based on a large number of repetitions, the bootstrap quantile $q_{\boldsymbol{v},1-\alpha}^{\star}$ is determined.
@@
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{\star}&=11\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{\star}\right),\\ &\varphi_{\boldsymbol{r}}^{\star}&=11\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{\star}\right).\end{array}$
For the investigation of hypotheses regarding correlations, even one further approach exists to obtain the required quantiles. It is called a Taylor-based Monte-Carlo technique and can be seen as a mixture between Monte-Carlo and bootstrap, using a refined version of the expansion in equation (2) to simulate the asymptotic distribution directly. Further details are provided in Sattler and Pauly (2023).
@@
Despite its high flexibility, the linear hypotheses introduced in (1) did not allow the formulation of all structures for covariance and correlation matrices. In particular, the autoregressive structure and all of their versions include some proportionality between the matrix components that cannot be captured through linear constraints on the vectorized covariance or correlation matrices. To accommodate such nonlinear structures, we consider transformations of the parameter vectors. Therefore, a kind of differentiable transformation $\boldsymbol{f}$ of the vectorized covariance resp. $\boldsymbol{g}$ of the correlation is required. Therewith, the hypothesis can be formulated through $\mathcal{H}_{0}^{\boldsymbol{v},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{v}}\boldsymbol{f}(\boldsymbol{v})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{v}}$ or $\mathcal{H}_{0}^{\boldsymbol{r},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{r}}\boldsymbol{g}(\boldsymbol{r})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{r}}$. By replacing the vectorized covariance resp. correlation by their transformed pendant, and adding the corresponding Jacobian matrix, the above-mentioned tests can be adapted. A detailed discussion, along with a useful transformation, referred to as the subdiagonal-mean-ratio, is provided in Sattler and Dobler (2024).
Although these tests are very versatile through the less restrictive assumptions and the general class of hypotheses, they do not by themselves yield insight into the specific reason a hypothesis is rejected — e.g., whether the differences arise due to variance differences, dependence structure, or both. Since correlation matrices reflect the dependency structure but contain no information about the components’ variances, but covariance matrices contain no standardized dependency measure, both on their own are not the best solution. To address this issue, Sattler and Pauly (2023) proposes a combined test for simultaneously assessing group differences in both covariance and correlation. This approach is based on the concept of Multiple Contrast Testing (see, e.g., Bretz et al. (2001)) and is applicable when comparing two groups ($a=2$). In case of rejection of equal covariance matrices, it also provides information on whether the dependency, the variance, or both are responsible for it. This test is based on a vector-value test statistic

### v2

@@ Abstract
In addition to the commonly analysed measures of location, dispersion measurements such as variance and correlation provide many valuable information. Consequently, they play a crucial role in multivariate statistics, which leads to tests regarding covariance and correlation matrices. Furthermore, also the structure of these matrices leads to important hypotheses of interest, since it contains substantial information about the underlying model. In fact, assumptions regarding the structures of covariance and correlation matrices are often fundamental in statistical modelling and testing. In this context, semi-parametric settings with minimal distributional assumptions and very general hypotheses are essential for enabling manifold usage. The freely available package CovCorTest provides suitable tests addressing all issues mentioned above, using bootstrap and similar techniques to achieve good performance, particularly in small samples. Additionally, the package offers flexible specification options for the hypotheses under investigation in two central tests, accommodating users with varying levels of expertise, which results in high flexibility and user-friendliness at the same time. This paper also presents the application of CovCorTest for various issues, illustrated by multiple examples, where the tests are applied to a real-world dataset.
@@ Introduction
Tests for measures of location are widely used in the analysis of multivariate data, such as those arising from repeated measures designs or multiple measurements on the same subject. While expectation vectors are a classical choice for such analyses (Konietschke et al. (2020)), alternative parameters such as quantiles (Baumeister et al. (2024)), relative effects (Rubarth et al. (2022) or Dobler et al. (2020)) and others have been proposed. In any case, measures of dispersion contain substantial information and are, therefore, of central importance in multivariate settings. Especially, covariance and correlation matrices are excellent parameters for analyses, which is, among other reasons, because:
@@
  • In statistical modelling, covariance and correlation matrices and their structures (e.g. autoregressive or compound symmetry) are essential for understanding and describing the underlying processes.
@@
Due to their importance, a variety of statistical procedures have been developed to test hypotheses concerning covariance matrices. Unfortunately, most existing methods rely on restrictive distributional assumptions, such as normality in Box (1953) and Zhong et al. (2017), or specific conditions on the characteristic function in Gupta and Xu (2006), which limit their applicability. Furthermore, several approaches are limited to specific settings, for example, two group settings as in Sakaori (2002) or Omelka and Pauly (2012). Moreover, they typically allow the investigation of only one or two concrete hypotheses, such as equality of covariance matrices (e.g., Yang and DeGruttola (2012) or Zhu et al. (2002)) or equality of two correlation matrices (e.g., Jennrich (1970)). This also holds for the software packages implementing these procedures, such as the R-packages psych(Revelle (2019)) or CovTestR(Barnard and Young (2018)).
@@
In contrast, the framework introduced in Sattler et al. (2022) allows a broad class of hypotheses, building on methodologies familiar from the analysis of location parameters. At the same time, the less restrictive semi-parametric setting used therein allows broad applicability. Thereto, test statistics are constructed as quadratic forms and Monte-Carlo techniques and resampling approaches are used to obtain the required quantiles and achieve good small sample performance. This approach was further expanded in Sattler and Pauly (2023) to quite general hypotheses regarding correlation matrices, where the focus was on the Anova-type statistic, and additionally, a Taylor-based Monte-Carlo approach was incorporated. The corresponding R package CovCorTest (Sattler et al. (2025)) includes the adequate tests from both papers to provide consistent procedures for investigating covariance and correlation matrices, and moreover also combined tests for equality of covariance and correlation matrices at the same time. Additionally, specific structures of covariance and correlation matrices, such as the Toeplitz form, are of great interest, as they provide meaningful insights into the underlying model. Besides, specific structures are usual model assumptions, but tests regarding these are often insufficient. Most existing approaches are limited to specific structures such as compound symmetry (Winer et al. (1991) or Votaw (1948)) or first-order autoregressive (McKeown and Johnson (1996)) or focus exclusively on either covariance or correlation matrices (Steiger (1980)). Conversely, the recent work by Sattler and Dobler (2024) introduces a unified framework for testing a large class of structural hypotheses, known as the linear covariance model, respectively the linear correlation model, and even further structures. These tests, which are related to the approaches presented in Sattler et al. (2022) and Sattler and Pauly (2023), have also been integrated into the package to enable comprehensive and coherent testing of various aspects of covariance and correlation matrices in one place.
@@
test_correlation_structure and test_combined), which are as analogous as possible to make it more user-friendly. This uniform design allows researchers to apply the functions in analogous situations, targeting different aspects of the data. Moreover, this approach allows for the investigation of all related hypotheses within a single package, while the coherent construction and methodology enhance the comparability of the results. Furthermore, to support users with varying levels of experience, the covariance and correlation tests allow not only the selection of predefined hypotheses, but also the specification of custom hypotheses by providing a hypothesis matrix and the corresponding vector. Since for the remaining tests, either only one hypothesis is possible (test_combined) or only predefined structures can be tested, no such differentiation is necessary there.
Because all underlying approaches and tests are based on empirical covariance and correlation matrices, our functions should only be used for metric data, although further assumptions, such as a repeated-measures design, are not required.
In the subsequent section, we will introduce the statistical multivariate model, the investigated null hypotheses, and the underlying assumptions. Afterwards, we present the implemented inference procedures, followed by the applications from the R package CovCorTest illustrated by some examples using a real-world dataset. After all, a conclusion and outlook complete the paper.
@@
We consider a general multigroup setting with $a\in{\mathbb{N}}$ groups, where each group i=1,,ai=1,...,a contains $n_{i}\in{\mathbb{N}}$ observations. The total number of observations is denoted by $N=\sum_{i=1}^{a}n_{i}$, allowing for potentially unbalanced designs in which group sizes may differ.
@@
where $\boldsymbol{\mu}_{i}\in R^{d}$ represents the group-specific mean vector, and $\boldsymbol{\epsilon}_{ik}\in R^{d}$ denotes the random deviation for subject $k=1,...,n_{i}$ from the $i$-th group. Here, splitting the indices allows considering factorial designs and longitudinal data, as demonstrated, e.g., in Pauly et al. (2015). We assume the residuals $\boldsymbol{\epsilon}_{i1},\dots,\boldsymbol{\epsilon}_{in_{i}}$ to be centred $\mathop{\operator@font I\hskip-1.5ptE\hskip 0.13pt}(\boldsymbol{\epsilon}_{i1})={\bf 0}_{d}$ and i.i.d. within each group, while across different groups only independence is needed, but differing distributions are possible. Further conditions for i=1,,ai=1,...,a are
@@
    $\operatorname{Var}(\boldsymbol{\epsilon}_{i1\ell})=V_{i\ell\ell}>0$,  $\ell=1,...,d$
@@
Since both covariance and correlation matrices are symmetric, it is natural to consider an appropriate vectorisation to transform these matrices into vectors containing the same information as the original matrix. For the covariance matrices, the so-called half-vectorisation operation $\operatorname{vech}$ is used, which extracts the upper-triangular (including diagonal) elements of a $d\times d$ matrix, yielding a $p=d(d+1)/2$ dimensional vector. The upper-half-vectorisation $\operatorname{vech}^{-}$, on the contrary, only takes the values above the diagonal into a $p_{-}=d(d-1)/2$ dimensional vector, which is useful for correlation matrices whose diagonal entries are always one and hence contain no information. Now, based on hypothesis matrices $\boldsymbol{C}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}\times p}$ respectively $\boldsymbol{C}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}\times p_{-}}$ and vectors $\boldsymbol{\zeta}_{\boldsymbol{v}}\in{\mathbb{R}}^{m_{1}}$, $\boldsymbol{\zeta}_{\boldsymbol{r}}\in{\mathbb{R}}^{m_{2}}$ general linear hypotheses can be formulated. These hypotheses are defined based on the pooled covariance vector $\boldsymbol{v}=(\operatorname{vech}(\boldsymbol{V}_{1})^{\top},\dots,\operatorname{vech}(\boldsymbol{V}_{a})^{\top})^{\top}$ respectively the pooled correlation vector $\boldsymbol{r}=(\operatorname{vech}^{-}(\boldsymbol{R}_{1})^{\top},\dots,\operatorname{vech}^{-}(\boldsymbol{R}_{a})^{\top})^{\top}$ as
@@
For correlation matrices, analogous—but typically fewer—structures exist, which often arise as heterogeneous versions of other structures such as heterogeneous autoregressive, heterogeneous compound symmetry or heterogeneous Toeplitz.
Except for the autoregressive structure and all its versions, all of these covariance and correlation structures can be expressed within the general hypothesis framework using appropriate choices of $\boldsymbol{C}$ and $\boldsymbol{\zeta}$, as shown in Sattler and Dobler (2024). Therein they introduced the linear covariance structure model, comparable to Anderson (1973) or Szatrowski (1980), given for $q<p$ through
$\mathcal{V}=\{\boldsymbol{V}(\boldsymbol{\theta})\in R^{d\times d}:\boldsymbol{V}(\boldsymbol{\theta})=\boldsymbol{V}_{0}+\theta_{1}\boldsymbol{V}_{1}+...+\theta_{q}\boldsymbol{V}_{q},\ \boldsymbol{\theta}=(\theta_{1},\dots,\theta_{q})^{\top}\in R^{q}\}\cap\mathcal{COV}_{d\times d}$ (2)
where $\boldsymbol{V}_{0},...,\boldsymbol{V}_{q}$ are known linearly independent symmetric matrices with $\boldsymbol{V}_{1},...,\boldsymbol{V}_{q}\neq{\bf 0}_{d\times d}$ and $\mathcal{COV}_{d\times d}\subset R^{d\times d}$ is the cone of covariance matrices, i.e., symmetric positive semi-definite matrices. In an analogous way, also the linear correlation structure model was defined. A central result thereof is that each covariance matrix structure from one of these models can be expressed through $\mathcal{H}_{0}^{\boldsymbol{v}}:\boldsymbol{C}_{\boldsymbol{v}}\boldsymbol{v}=\boldsymbol{\zeta}_{\boldsymbol{v}}$ respectively $\mathcal{H}_{0}^{\boldsymbol{r}}:\boldsymbol{C}_{\boldsymbol{r}}\boldsymbol{r}=\boldsymbol{\zeta}_{\boldsymbol{r}}$. Although these expressions are not unique, the package includes a function get_hypothesis which for given $\operatorname{vech}(\boldsymbol{V}_{0}),(\operatorname{vech}(\boldsymbol{V}_{1}),...,\operatorname{vech}(\boldsymbol{V}_{q}))$ resp. $\operatorname{vech}^{-}(\boldsymbol{R}_{0}),(\operatorname{vech}^{-}(\boldsymbol{R}_{1}),...,\operatorname{vech}^{-}(\boldsymbol{R}_{q}))$ (for the linear correlation structure model) returns a possible hypothesis matrix $\boldsymbol{C}$ and corresponding vector $\boldsymbol{\zeta}$.
This highlights the versatility and generality of the presented model in capturing also a wide range of relevant covariance and correlation structures.
@@
To evaluate the hypotheses formulated in the previous section, we develop inference procedures based on the asymptotic distribution of vectorised covariance and correlation estimates.
For our hypotheses regarding covariance matrices, we consider the vectorised empirical group covariances $\widehat{\boldsymbol{v}}_{i}:=\operatorname{vech}\left((n_{i}-1)^{-1}\sum_{i=1}^{n_{i}}(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})(\boldsymbol{X}_{ik}-\overline{\boldsymbol{X}}_{i})^{\top})\right)$ where $\overline{\boldsymbol{X}}_{i}$ is the sample mean of group $i$. Stacking these vectors across all $a$ groups yields the pooled vector $\widehat{\boldsymbol{v}}=(\widehat{\boldsymbol{v}}_{1}^{\top},...,\widehat{\boldsymbol{v}}_{a}^{\top})^{\top}$. Then, for developing a statistical test, we use that under assumption (A2) and (A3) for $i=1,\dots,a$, the group-wise covariance vectors satisfy
@@
In the same way, $\widehat{\boldsymbol{r}}_{i}$ denotes the empirical counterpart of the upper-half-vectorised correlation matrix of group $i$ and $\widehat{\boldsymbol{r}}$ is the corresponding pooled vector. Following the results in Sattler and Pauly (2023) (where also the exact form can be found), there exists a continuous matrix-valued function $\boldsymbol{M}\in{\mathbb{R}}^{p_{-}\times p}$ such that
[…]
@@
There also exist alternative versions of the ATS, which, for example, allow the approximation through an F-distribution (see e.g. Brunner et al. (2019) or Thiel et al. (2024)). However, for the statistics defined in (4) and (5), the asymptotic null distribution is known, enabling valid inference without stringent distributional assumptions.
@@
A Monte-Carlo technique is one useful approach to obtain the required quantiles of these weighted sums for this statistic. For the covariance matrix, with our consistent covariance estimator $\widehat{\boldsymbol{\Sigma}}$, we get estimators $\hat{\lambda}_{1},...,\hat{\lambda}_{m_{1}}$ for the unknown eigenvalues. With these weights we generate appropriate realisations $B_{1},...,B_{m_{2}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\chi_{1}^{2}$. These realisations and our estimators allow calculating a weighted sum $\sum\limits_{\ell=1}^{m_{2}}\hat{\lambda}_{\ell}B_{\ell}$, which converges in distribution against the above limit distribution. Repeating this procedure many times allows to calculate the required Monte-Carlo quantiles, and with an analogous procedure for the correlation for $\alpha\in(0,1)$, we get asymptotic level $\alpha$ tests through
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{MC}&=1\hskip-3.01385pt1\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{MC}\right),\\ &\varphi_{\boldsymbol{r}}^{MC}&=1\hskip-3.01385pt1\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{MC}\right).\end{array}$
Another technique for obtaining quantiles that requires more computational effort but usually yields better small-sample performance is a bootstrap approach. For hypotheses regarding covariance matrices, the parametric bootstrap mimics the limit distribution and afterwards the structure of the test statistic by generating bootstrap vectors $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}_{p}\left({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i}\right),$ based on given realisations $\boldsymbol{X}_{i1},...,\boldsymbol{X}_{in_{i}}$ with estimators $\widehat{\boldsymbol{\Sigma}}_{i}$. Then $\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$, the empirical covariance matrix of the bootstrap sample $\boldsymbol{Z}_{i1}^{\star},...,\boldsymbol{Z}_{in_{i}}^{\star}$ and $\widehat{\boldsymbol{\Sigma}}^{\star}:=\bigoplus_{i=1}^{a}{N}/{n_{i}}\cdot\widehat{\boldsymbol{\Sigma}}_{i}^{\star}$ are calculated. Following the bootstrap realization $ATS_{\boldsymbol{v}}^{\star}=N[\boldsymbol{C}\overline{\boldsymbol{Z}}^{\star}]^{\top}[\boldsymbol{C}\ \overline{\boldsymbol{Z}}^{\star}]\big/\operatorname{tr}(\boldsymbol{C}\widehat{\boldsymbol{\Sigma}}^{\star}\boldsymbol{C}^{\top})$ is calculated, using the bootstrap means $\overline{\boldsymbol{Z}}_{i}^{\star}=n_{i}^{-1}\sum_{k=1}^{n_{i}}\boldsymbol{Z}_{i}k^{\star}$ and $\overline{\boldsymbol{Z}}^{\star}=({\overline{\boldsymbol{Z}}_{1}^{\star}}^{\top},...,{\overline{\boldsymbol{Z}}_{a}^{\star}}^{\top})^{\top}$. Then, based on a large number of repetitions, the bootstrap quantile $q_{\boldsymbol{v},1-\alpha}^{\star}$ is determined.
@@
$\begin{array}[]{lrl}&\varphi_{\boldsymbol{v}}^{\star}&=1\hskip-3.01385pt1\left(ATS_{\boldsymbol{v}}>q_{\boldsymbol{v},1-\alpha}^{\star}\right),\\ &\varphi_{\boldsymbol{r}}^{\star}&=1\hskip-3.01385pt1\left(ATS_{\boldsymbol{r}}>q_{\boldsymbol{r},1-\alpha}^{\star}\right).\end{array}$
For the investigation of hypotheses regarding correlations, an additional approach exists to obtain the required quantiles. It is called a Taylor-based Monte-Carlo technique and can be seen as a mixture between Monte-Carlo and bootstrap, using a refined version of the expansion in equation (3) to simulate the asymptotic distribution directly. Further details are provided in Sattler and Pauly (2023).
@@
Despite its high flexibility, the linear hypotheses introduced in (1) did not allow the formulation of all structures for covariance and correlation matrices. In particular, the autoregressive structure and all of its versions include some proportionality between the matrix components that cannot be captured through linear constraints on the vectorised covariance or correlation matrices. To accommodate such nonlinear structures, we consider transformations of the parameter vectors. Therefore, a kind of differentiable transformation $\boldsymbol{f}$ of the vectorised covariance resp. $\boldsymbol{g}$ of the correlation is required. Therewith, the hypothesis can be formulated through $\mathcal{H}_{0}^{\boldsymbol{v},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{v}}\boldsymbol{f}(\boldsymbol{v})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{v}}$ or $\mathcal{H}_{0}^{\boldsymbol{r},s}:\widetilde{\boldsymbol{C}}_{\boldsymbol{r}}\boldsymbol{g}(\boldsymbol{r})=\widetilde{\boldsymbol{\zeta}}_{\boldsymbol{r}}$. By replacing the vectorised covariance matrix respectively correlation matrix by their transformed pendant, and adding the corresponding Jacobian matrix, the above-mentioned tests can be adapted. A detailed discussion, along with a useful transformation, referred to as the subdiagonal-mean-ratio, is provided in Sattler and Dobler (2024).
Although these tests are very versatile through the less restrictive assumptions and the general class of hypotheses, they do not by themselves yield insight into the specific reason a hypothesis is rejected — e.g., whether the differences arise due to variance differences, dependence structure, or both. Since correlation matrices reflect the dependency structure but contain no information about the components’ variances, and covariance matrices contain no standardised dependency measure, neither is the best solution on its own. To address this issue, Sattler and Pauly (2023) proposes a combined test for simultaneously assessing group differences in both covariance and correlation. This approach is based on the concept of Multiple Contrast Testing (see, e.g., Bretz et al. (2001)) and is applicable when comparing two groups ($a=2$). In case of rejection of equal covariance matrices, it also provides information on whether the dependency, the variance, or both are responsible for it. This test is based on a vector-value test statistic